# Constraining Power of Wavelet vs. Power Spectrum Statistics for CMB Lensing and Weak Lensing with Learned Binning

### v1

@@ 3.3 Wavelet Phase Harmonics
As in WST, $\psi_{j,l}$ is a localized wavelet characterized by a scale parameter $j$ and a directional parameter $l$. In general, $A_{j,l}$ will be complex valued. We apply a phase harmonic to this, defined as:
@@
These statistics are chosen following Gatti et al. (2023), with the only change being that we allow for more opening angles in $C_{01}$ and $C_{10}$ statistics. For simplicity, we only consider the real component of the final statistics.
In this work we use the package PyWPH (Regaldo-Saint Blancard et al., 2021) to apply WPH and use the bump-steerable wavelets, which are given by the following mother wavelet:
@@
The equation above is for Fourier wavenumbers in the range $0<k<2\xi_{0}$ and $|\text{arg}(k)|<\pi/2$; for all other values $\psi=0$. The normalization and the central frequency $\xi_{0}=0.85\pi$ are set as in Regaldo-Saint Blancard et al. (2021). All other wavelets are generated by dilating or rotating the mother wavelet. In real space:
@@
4 Analysis
We use the Fisher information approach to compare the constraining power of our different summary statistics. In doing so, we are assuming that our summary statistics are Gaussian distributed, which we test and confirm in Appendix D. To fairly compare different summary statistics, we compress them with a novel learned binning approach. This approach compresses the statistics to a set number of bins while maximizing the determinant of the Fisher matrix.
4.1 Fisher Forecast
@@
To unbias our inverse covariance, we apply the Hartlap correction (Hartlap et al., 2006); however, this correction alone does not result in an unbiased Fisher matrix. Noise in the estimates of the derivatives causes the Fisher matrix to systematically overestimate the constraining power of a summary statistic (Coulton & Wandelt, 2023).
To address this, Coulton & Wandelt (2023) introduce a compressed Fisher matrix, $F^{\text{comp}}$. This Fisher matrix is constructed as in Eq. 24 but using summary statistics that have been compressed with the MOPED (Heavens et al., 2000) algorithm. Perfect compression of the summary statistics would lead to no information loss, but the compression relies on the matrices $D$ and $C$. Coulton & Wandelt (2023) show that $F^{\text{comp}}$ will systematically underestimate the constraining power if $D$ and $C$ are estimated from simulations via Monte Carlo methods before compressing the statistics. Taking the geometric mean of $F$ and $F^{\text{comp}}$ gives the combined Fisher matrix, $F^{\text{Combined}}$, in which the opposing biases of the two inputs cancel, yielding unbiased constraints. It is this combined Fisher matrix and the resulting parameter covariance matrix, $(F^{\text{Combined}})^{-1}$, which we use for all shown constraints.
@@
Before constructing the combined Fisher matrix, we need to compress our summary statistics in order to get an invertible data covariance matrix. Compression is necessary since the finite number of realizations used to evaluate the covariance matrix places an upper bound on its rank. Since all of our summary statistics have scale dependence, in this work we implement a novel learned binning (over scale) approach to compress our statistics, while retaining interpretability. We explain our method below.
4.2 Learned Binning Compression
@@
Figure 4: The determinant (normalized to have a maximum of 11) of the Fisher matrix from Eq. 24 over pivot space for Planck $\kappa_{\text{CMB}}$ $C_{\ell}$’s. Only one pivot is used to allow for visualization via a heatmap, so the axes denote the $\ell$ and bin position of the one pivot.
In this section, the initial summary statistic will be $M$ dimensional and it will be compressed to $N$ dimensions, with $N<M$. Any binning scheme can be viewed as a particular linear transformation within $\mathbb{R}^{N\times M}$. We will refer to the general binning matrix as $B$, and to compress a data vector $\vec{d}$ we perform $B\vec{d}$.
@@
In this way, we are now inverting an $N\times N$ matrix instead of an $M\times M$ matrix. Therefore, we want the value of $N$ to be much smaller than the number of fiducial simulations. Everything so far simply applies to general $N\times M$ linear transformations beyond just binning matrices. Compared to general linear transformations, binning is more intuitive and interpretable while having fewer free parameters, thus reducing the chance of overfitting.
To enable low-dimensional parameterizations of binning matrices, we use a function $f(x):[1,M]\rightarrow[1,N]$ to specify the positions of nonzero matrix entries. Specifically, the $m^{\text{th}}$ column of the binning matrix has a single nonzero entry equal to $1$ at the integer index closest to $f(m)$. In this work, $f(x)$ is a linear interpolation between $(1,1)$, learned pivots, and $(M,N)$. Further details on the conditions that $f(x)$ must satisfy can be found in Appendix C. An example of such a construction is shown in Figure 3. We will often refer to the summary statistic index and bin number of the pivot as the “xx” or “yy” components of the pivot.
To compress summary statistics, we first pick a number of pivots for the binning matrix. Some summary statistics (cross-$C_{\ell}$’s, WST, and WPH, in this work) can be thought of as consisting of multiple subclasses of statistics that depend exclusively on scale. Taking WST as an example, $S_{1}$ coefficients and $S_{2}$ coefficients cannot be naturally compared based on scale. Additionally, within the $S_{2}$ coefficients, statistics with different opening angles cannot be naturally ordered based on scale. Therefore, for WST we split the coefficients into $5$ subclasses that have only scale dependence ($S_{1}$ and the four different opening angle components of $S_{2}$). These statistics are individually ordered based on scale and then appended together to form our full ordered summary statistic. We fix the $x$ coordinates of certain pivots to the transitions between subclasses. This effectively introduces parameters which, in combination with the total number of bins, exclusively control how many bins are allocated to each subclass.
@@
Each point in pivot space corresponds to a binning of the summary statistics which can be used to create a unique Fisher matrix. To retain as much information as possible, we attempt to maximize $\det(F)$. For the sake of efficiency, this optimization uses the Fisher matrix from Eq. 24 instead of the combined Fisher matrix. The relatively low number of free parameters ($\mathcal{O}(\sim 10)$ in this work) reduces the probability of overfitting compared to a general linear compression with $N\times M$ free parameters.
@@
When only using one pivot, the pivot space is two dimensional, and $\det(F)$ can be plotted as a heatmap. An example of this for Planck $\kappa_{\text{CMB}}$ $C_{\ell}$’s is shown in Figure 4. This plot shows that $\det(F)$ has a genuine signal as a function over pivot space, but that there is also numerical noise that comes from having a finite number of simulations. Since we maximize $\det(F)$, we could end up fitting to a local maximum in the noise, which would lead to us overestimating the strength of constraints.
@@
To avoid overfitting, we randomly split our fiducial and derivative simulations in half, learn pivot positions on one half of the data, and then generate contours based on the second half of the data. While we use the Fisher matrix from Eq. 24 to learn how to bin our summary statistics, once compressed, we use the combined Fisher matrix to create contour plots. We repeat this process for $1000$ different splits and get a covariance matrix for our cosmological parameters for each split. We take our final covariance matrix to be the average of these matrices.
@@
We use $15$ bins since we find that this choice yields binned summary statistics which follow a Gaussian distribution (see Appendix D). With a total of $15$ bins, it would be feasible to attempt to find the size of every bin individually. However, this approach would struggle to generalize to future works if more bins were used. Determining the optimal placements of bin boundaries for $N$ bins consists of maximizing $\det(F)$ over an N1N-1 dimensional space. When using our pivots approach, the dimension of the pivot space $\det(F)$ is maximized over is completely decoupled from $N$. This allows one to control the number of free parameters.
@@
In general, we note that the compression of summary statistics is a non-trivial problem, and a variety of methods have been proposed in the literature (Park et al., 2025). Many of these techniques, however, require training on simulations at varying cosmologies, which is not feasible in our setting where only a fiducial covariance matrix and its derivatives are available. Other possible alternatives in this regime are Principal Component Analysis (PCA), finding ideal subsets of observables (Szewciw et al., 2022), and the more widely used MOPED algorithm (Heavens et al., 2000). In principle, MOPED would be the most principled choice, as it is designed to maximize the Fisher information. In practice, however, its performance is limited in our case: it relies on stable estimates of derivatives and on the inversion of a high-dimensional covariance matrix, both of which can be noisy when only a modest number of simulations are available. This issue is exacerbated by the need to split the simulations into independent subsets for training and validation to avoid overfitting, which further reduces the effective number available. Although our learned-binning scheme also maximizes Fisher information, the compression matrix is constrained to a simple, nearly diagonal form that depends on only a handful of pivot parameters. This structure makes it substantially less sensitive to noise than the dense compression vectors produced by MOPED, while preserving interpretability. The additional use of cross-validation further mitigates overfitting to fluctuations in the Fisher matrix, yielding a robust and practical method.
5 Results
@@
We now turn to forecasted constraints on $\Omega_{m}$ and $\sigma_{8}$ after binning our summary statistics. We first compare auto $C_{\ell}$’s and WST applied to $\kappa_{\text{CMB}}$ maps in Section 5.1. Next, we look at cross-$C_{\ell}$’s and WPH applied to $\kappa_{\text{CMB}}\times\kappa_{\text{WL}}$ maps in Section 5.2. We note that results for WST and WPH in the next section are not combined with those from the power spectrum.
5.1 CMB Lensing
@@
Figure 5: We show $1\sigma$ and $2\sigma$ constraints on $\Omega_{m}$ and $\sigma_{8}$ from Planck $\kappa_{\text{CMB}}$, using the $C_{\ell}$’s (blue) and the WST coefficients (orange). These constraints are obtained using our learned-binning approach. The WST improves constraints by a factor of $1.3$ for $\Omega_{m}$ and a factor of $1.7$ for $\sigma_{8}$.
@@
For WST, we do not use dyadic sequencing; instead we have $j$ from Eq. 11 take $50$ different values from 0 to $6$. This allows for more flexibility in the learned binning by increasing the size of the initial data vector. For angular spacing, we pick $L=4$. With these choices, there are 5150 WST coefficients. For $C_{\ell}$’s, our resolution limits the maximum $\ell$ to 3071, so after discarding the dipole we have 3070 $C_{\ell}$ values.
For the learned binning, ordering the original summary statistic is important. Entries that are near each other will likely end up being binned together, so we want these entries to be highly correlated to ensure minimal loss of information. For $C_{\ell}$’s this is done by ordering based on $\ell$. $S_{1}$ coefficients are likewise ordered based on $j$. For $S_{2}$ coefficients, we have two scales and an opening angle, making an ordering scheme more arbitrary. In this work, we divide the $S_{2}$ coefficients into $4$ different subclasses based on the opening angle. In each subclass, we order based on $j_{2}$. When multiple coefficients have the same opening angle and $j_{2}$ value, we order based on $j_{1}$.
For the $C_{\ell}$’s, we find that allowing for one pivot can slightly improve constraints, but going to two gives little additional gain. Therefore, for $C_{\ell}$’s we search for one pivot. For WST, we lock four pivots to the transitions from $S_{1}$ to $S_{2}$ and between the different opening angles of $S_{2}$. Since one dimension of the pivot is locked, this has four degrees of freedom. We find improvements by adding one more arbitrary pivot that can be fit anywhere, leading to a $6$ dimensional search space for the pivots, with no significant improvements going beyond this.
@@
The relative performance across surveys reflects the interplay between noise levels and sky coverage. Planck has the highest noise but covers nearly the full sky, providing a vast number of modes; in this regime, the power spectrum alone is less constraining, allowing WST to recover additional information from non-Gaussian features. ACT lies in an intermediate regime, with lower noise but smaller area, where WST still improves constraints by capturing complementary statistics. SO and SPT deliver the lowest-noise reconstructions; in this limit, most of the cosmological information is already encoded in the power spectrum, leaving limited additional gain from WST.
5.2 CMB Lensing X Weak Lensing
Figure 6: Constraints for Planck $\times$ Euclid DR2 ($\kappa_{\text{CMB}}\times\kappa_{\text{WL}}$) from WPH. The learned binning (orange) outperforms the initial binning with no pivots (blue). Constraints are improved by factors of $2.7$ for $\Omega_{m}$ and $2.3$ for $\sigma_{8}$.
Figure 7: Constraints for Planck $\times$ Euclid DR2 ($\kappa_{\text{CMB}}\times\kappa_{\text{WL}}$) from cross-$C_{\ell}$’s and WPH using our learned-binning approach. The $\Omega_{m}$ and $\sigma_{8}$ constraints are improved by factors of $2.7$ and $2.4$, respectively, when going from cross-$C_{\ell}$’s to WPH.

### v2

@@ III.3 Wavelet Phase Harmonics
As in WST, $\psi_{j,l}$ is a localized wavelet characterized by a scale parameter $j$ and a directional parameter $l$. In general, $A_{j,l}$ will be complex-valued. We apply a phase harmonic to this, defined as:
@@
These statistics are chosen following Gatti et al. [18], with the only change being that we allow for more opening angles in $C_{01}$ and $C_{10}$ statistics. For simplicity, we only consider the real component of the final statistics.
In this work, we use the package PyWPH [40] to apply WPH and use the bump-steerable wavelets, which are given by the following mother wavelet:
@@
The equation above is for Fourier wavenumbers in the range $0<k<2\xi_{0}$ and $|\text{arg}(k)|<\pi/2$; for all other values $\psi=0$. The normalization and the central frequency $\xi_{0}=0.85\pi$ are set as in Regaldo-Saint Blancard et al. [40]. All other wavelets are generated by dilating or rotating the mother wavelet. In real space:
@@
IV Analysis
We use the Fisher information approach to compare the constraining power of our different summary statistics. In doing so, we assume that our summary statistics are Gaussian-distributed, which we test and confirm in Appendix E. To fairly compare different summary statistics, we compress them with a novel learned binning approach. This approach compresses the statistics to a set number of bins while maximizing the determinant of the Fisher matrix.
IV.1 Fisher Forecast
@@
To unbias our inverse covariance, we apply the Hartlap correction [21]; however, this correction alone does not yield an unbiased Fisher matrix. Noise in derivative estimates causes the Fisher matrix to systematically overestimate the constraining power of a summary statistic [15].
To address this, Coulton and Wandelt [15] introduce a compressed Fisher matrix, $F^{\text{comp}}$. This Fisher matrix is constructed as in Eq. 24, but using summary statistics compressed with the MOPED [22] algorithm. Perfect compression of the summary statistics would lead to no information loss, but the compression relies on the matrices $D$ and $C$. Coulton and Wandelt [15] show that $F^{\text{comp}}$ will systematically underestimate the constraining power if $D$ and $C$ are estimated from simulations via Monte Carlo methods before compressing the statistics. Taking the geometric mean of $F$ and $F^{\text{comp}}$ gives the combined Fisher matrix, $F^{\text{Combined}}$, in which the opposing biases of the two inputs cancel, yielding unbiased constraints. It is this combined Fisher matrix and the resulting parameter covariance matrix, $(F^{\text{Combined}})^{-1}$, which we use for all shown constraints.
@@
Before constructing the combined Fisher matrix, we need to compress our summary statistics to obtain an invertible data covariance matrix. Compression is necessary since the finite number of realizations used to evaluate the covariance matrix places an upper bound on its rank. Since all of our summary statistics have scale dependence, in this work we implement a novel learned binning (over scale) approach to compress our statistics, while retaining interpretability. We explain our method below.
IV.2 Learned Binning Compression
@@
Figure 4: The determinant (normalized to have a maximum of 11) of the Fisher matrix from Eq. 24 over pivot space for Planck $\kappa_{\text{CMB}}$ $C_{\ell}$’s. Only one pivot is used to enable visualization via a heatmap, so the axes denote the $\ell$ and bin positions of the pivot.
In this section, the initial summary statistic will be $M$-dimensional, and it will be compressed to $N$ dimensions, with $N<M$. Any binning scheme can be viewed as a particular linear transformation within $\mathbb{R}^{N\times M}$. We will refer to the general binning matrix as $B$. To compress a data vector $\vec{d}$, we perform $B\vec{d}$.
@@
In this way, we are now inverting an $N\times N$ matrix instead of an $M\times M$ matrix. Therefore, we want the value of $N$ to be much smaller than the number of fiducial simulations. Everything so far simply applies to general $N\times M$ linear transformations beyond just binning matrices. Compared with general linear transformations, binning is more intuitive and interpretable while having fewer free parameters, thereby reducing the risk of overfitting.
To enable low-dimensional parameterizations of binning matrices, we use a function $f(x):[1,M]\rightarrow[1,N]$ to specify the positions of nonzero matrix entries. Specifically, the $m^{\text{th}}$ column of the binning matrix has a single nonzero entry equal to $1$ at the integer index closest to $f(m)$. In this work, $f(x)$ is a linear interpolation between $(1,1)$, learned pivots, and $(M,N)$. Further details on the conditions that $f(x)$ must satisfy can be found in Appendix D. An example of such a construction is shown in Figure 3. We will often refer to the summary statistic index and bin number of the pivot as the “xx” or “yy” components of the pivot.
To compress summary statistics, we first pick a number of pivots for the binning matrix. Some summary statistics (cross-$C_{\ell}$’s, WST, and WPH, in this work) can be thought of as consisting of multiple subclasses of statistics that depend exclusively on scale. Taking WST as an example, $S_{1}$ and $S_{2}$ coefficients cannot be naturally compared based on scale. Additionally, within the $S_{2}$ coefficients, statistics with different opening angles cannot be naturally ordered based on scale. Therefore, for WST, we split the coefficients into $5$ subclasses that have only scale dependence ($S_{1}$ and the four different opening angle components of $S_{2}$). These statistics are ordered individually by scale and then appended to form our full ordered summary statistic. We fix the $x$ coordinates of certain pivots to the transitions between subclasses. This effectively introduces parameters which, in combination with the total number of bins, exclusively control how many bins are allocated to each subclass.
@@
Each point in pivot space corresponds to a binning of the summary statistics which can be used to create a unique Fisher matrix. To retain as much information as possible, we attempt to maximize $\det(F)$. For the sake of efficiency, this optimization uses the Fisher matrix from Eq. 24 instead of the combined Fisher matrix. The relatively low number of free parameters ($\mathcal{O}(\sim 10)$ in this work) reduces the probability of overfitting compared to a general linear compression with $N\times M$ free parameters. An outline of our optimization algorithm is given in Appendix D.
@@
When only using one pivot, the pivot space is two-dimensional, and $\det(F)$ can be plotted as a heatmap. An example of this for Planck $\kappa_{\text{CMB}}$ $C_{\ell}$’s is shown in Figure 4. This plot shows that $\det(F)$ has a genuine signal as a function over pivot space, but that there is also numerical noise that comes from having a finite number of simulations. Since we maximize $\det(F)$, we could end up fitting to a local maximum in the noise, which would lead to us overestimating the strength of constraints.
@@
To avoid overfitting, we randomly split our fiducial and derivative simulations in half, learn pivot positions on one half, and then generate contours from the other half. While we use the Fisher matrix from Eq. 24 to learn how to bin our summary statistics, once compressed, we use the combined Fisher matrix to create contour plots. We repeat this process for $1000$ different splits and get a covariance matrix for our cosmological parameters for each split. We take our final covariance matrix to be the average of these matrices.
@@
We use $15$ bins since we find that this choice yields binned summary statistics that follow a Gaussian distribution (see Appendix E). With $15$ bins, it would be feasible to determine the size of each bin individually. However, this approach would struggle to generalize to future works if more bins were used. Determining the optimal placements of bin boundaries for $N$ bins consists of maximizing $\det(F)$ over an N1N-1 dimensional space. When using our pivot approach, the dimension of the pivot space, $\det(F)$, is maximized over and is completely decoupled from $N$. This allows one to control the number of free parameters.
@@
In general, we note that compressing summary statistics is a non-trivial problem, and a variety of methods have been proposed in the literature [35]. Many of these techniques, however, require training on simulations at varying cosmologies, which is not feasible in our setting, where only a fiducial covariance matrix and its derivatives are available. Other alternatives in this regime include Principal Component Analysis (PCA), finding optimal subsets of observables [45], and the widely used MOPED algorithm [22]. In principle, MOPED would be the most principled choice, as it is designed to maximize the Fisher information. In practice, however, its performance is limited in our case: it relies on stable estimates of derivatives and on inverting a high-dimensional covariance matrix, both of which can be noisy when only a modest number of simulations are available. This issue is exacerbated by the need to split the simulations into independent subsets for training and validation to avoid overfitting, which further reduces the effective number available. Although our learned-binning scheme also maximizes Fisher information, the compression matrix is constrained to a simple, nearly diagonal form that depends on only a handful of pivot parameters. This structure makes it substantially less sensitive to noise than the dense compression vectors produced by MOPED, while preserving interpretability. Using cross-validation further mitigates overfitting to fluctuations in the Fisher matrix, yielding a robust and practical method.
V Results
@@
We now turn to forecasted constraints on $\Omega_{m}$ and $\sigma_{8}$ after binning our summary statistics. We first compare auto $C_{\ell}$’s and WST applied to $\kappa_{\text{CMB}}$ maps in Section V.1. Next, we look at cross-$C_{\ell}$’s and WPH applied to $\kappa_{\text{CMB}}\times\kappa_{\text{WL}}$ maps in Section V.2. We note that results for WST and WPH in the next section are not combined with those from the power spectrum.
V.1 CMB Lensing
@@
Figure 5: We show $1\sigma$ and $2\sigma$ constraints on $\Omega_{m}$ and $\sigma_{8}$ from Planck $\kappa_{\text{CMB}}$, using the $C_{\ell}$’s (blue) and the WST coefficients (orange). These constraints are obtained using our learned-binning approach. The constraints are the same to within $5\%$.
@@
For WST, we do not use dyadic sequencing; instead, we have $j$ from Eq. 11 take $50$ different values from 0 to $6$. This allows greater flexibility in the learned binning by increasing the size of the initial data vector. For angular spacing, we pick $L=4$. With these choices, there are 5150 WST coefficients. For $C_{\ell}$’s, our resolution limits the maximum $\ell$ to 3071, so after discarding the dipole, we have 3070 $C_{\ell}$ values.
For the learned binning, ordering the original summary statistic is important. Entries that are near each other will likely be binned together, so we want these entries to be highly correlated to minimize information loss. For $C_{\ell}$’s, this is done by ordering based on $\ell$. $S_{1}$ coefficients are likewise ordered based on $j$. For $S_{2}$ coefficients, we have two scales and an opening angle, making the ordering scheme more arbitrary. In this work, we divide the $S_{2}$ coefficients into $4$ different subclasses based on the opening angle. In each subclass, we order based on $j_{2}$. When multiple coefficients have the same opening angle and $j_{2}$ value, we order based on $j_{1}$. Across surveys, we find that when looking at pivot positions, the $S_{1}$ coefficients are the most heavily weighted.
For the $C_{\ell}$’s, we find that allowing one pivot can slightly improve the constraints, but allowing two yields little additional gain. Therefore, for $C_{\ell}$’s we search for one pivot. For WST, we lock four pivots to the transitions from $S_{1}$ to $S_{2}$ and between the different opening angles of $S_{2}$. Since one dimension of the pivot is locked, this has four degrees of freedom. We find improvements by adding one additional arbitrary pivot that can be placed anywhere, resulting in a $6$-dimensional search space for the pivots, with no significant gains beyond this. Across surveys, we find that the first $\sim 1000$ $C_{\ell}$’s are consistently the most heavily weighted.
@@
When we look at the covariance of the two binned statistics, we find the expected positive correlation between $C_{\ell}$’s and $S_{1}$ coefficients. Additionally, the covariance matrix of the WST coefficients has significant off-diagonal entries, primarily between the $S_{2}$ coefficients. This highlights that while our learned binning algorithm can be useful in diagonalizing the covariance matrix, it is only possible if correlations are originally among entries close enough together to be in the same bin. This corresponds to an initial non-binned covariance matrix with significant entries only near the diagonal.
V.2 CMB Lensing X Weak Lensing
Figure 6: Constraints for ACT $\times$ Euclid DR2 ($\kappa_{\text{CMB}}\times\kappa_{\text{WL}}$) from WPH. The learned binning (orange) outperforms the initial binning with no pivots (blue). Constraints are improved by factors of $2.7$ for $\Omega_{m}$ and $2.5$ for $\sigma_{8}$.
Figure 7: Constraints for Planck $\times$ Euclid DR2 ($\kappa_{\text{CMB}}\times\kappa_{\text{WL}}$) from cross-$C_{\ell}$’s and WPH using our learned-binning approach. The $\Omega_{m}$ and $\sigma_{8}$ constraints are improved by factors of $2.3$ and $2.2$, respectively, when going from cross-$C_{\ell}$’s to WPH.